# Position: Many generalization measures for deep learning are fragile

### v1

@@ Abstract
A wide variety of generalization measures have been applied to deep neural networks (DNNs). Although obtaining tight bounds remains challenging, such measures are often assumed to reproduce qualitative generalization trends. In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks— are fragile: small training modifications that barely affect the underlying DNN can substantially change a measure’s value, trend, or scaling behaviour. For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants can reverse the slope of a learning curve in widely used generalisation measures like the path norm. We also identify subtler forms of fragility. For instance, the PAC-Bayes origin measure is regarded as one of the most reliable, and is indeed less sensitive to hyperparameter tweaks than many other measures. However, it completely fails to capture differences in data complexity across learning curves. This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behaviour, in learning curves, but which is not a post-mortem measure. Beyond demonstrating that many bounds - such as path, spectral and Frobenius norms, flatness proxies, and deterministic PAC–Bayes surrogates- are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
@@ 1 Introduction
Classical generalization theories—based on VC dimension, Rademacher complexity and related tools—have a long history in machine learning. They were developed with two intertwined aims: to furnish tight mathematical guarantees and to explain generalization. For deep networks the same aspiration has driven a profusion of bounds. While tight bounds have proven elusive, with some exceptions (Dziugaite & Roy, 2017; Pérez‐Ortiz et al., 2020; Valle-Pérez & Louis, 2020; Lotfi et al., 2022), the expectation has been that these measures can generate qualitative insight into the generalization behaviour of deep neural networks (DNNs). These bounds have typically focused on weights after training, which is why they are sometimes called post-mortem bounds. Large-scale comparisons have compared the behaviour of these measures, albeit often only at the sign level, e.g. does changing a setting produce the same a change in the underlying DNN and a generalization measure with the same sign (Jiang et al., 2019a; Dziugaite et al., 2020a).
@@
The widespread fragility we observe suggests that some generalisation measures may be showing correlation in some limited settings that may not be related to causation. Our position is therefore not only that many measures are fragile, but also that investigators should carefully audit their measures for fragility, using the kinds of tests we propose in this paper.
@@ 3 Training-hyperparameter fragility
In the left plot, we run SGD with momentum at $\mathrm{LR}=0.01$: both path‑norm curves start at a value of roughly ($\sim 10^{5}$) and decrease more or less log‑linearly with increasing sample size, with the margin‑normalized curve tracking about a decade lower. Both drop faster than the underlying ResNet‑50 test error. Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behaviour. A minor adjjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows. These widely different behaviours arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
The contrasting behaviour of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression. In that setting, for a fixed $\ell_{r}$ weight norm, qualitatively different scaling of the norm with sample size obtains depending on the $\ell_{p}$‑minimizer. More precisely, with $p\in[1,2]$, the $\ell_{r}$ weight norm is monotonic in $n$ if $r\leq 2(p-1)$ and plateaus if $r\leq 2(p-1)$  (Zhang & Louis, 2025). This implies that fixing which norm you measure does not fix its learning‑curve scaling, because the solution selected by optimization—through implicit (or explicit) bias, e.g., favoring different $\ell_{p}$‑minimizers—can induce different sample‑size scalings of that same norm. This change in qualitative behavior was also illustrated in diagonal linear networks where the implicit bias can be tuned by varying the learning dynamics (Zhang & Louis, 2025). For the more complex setup of path-norm for ResNet‑50 on FashinMNIST in Fig. 1, the optimizer and learning‑rate choices may analogously shift the deep model’s implicit bias, toggling the path‑norm scaling between monotone and U‑shaped, just as was seen for linear regression and deep linear networks.
@@ 5 Data-Complexity Fragility: Label Noise & Dataset Difficulty
Finally, the last panel in figure 3 illustrates how the function‑space ML‑PACBayes measure provides a non-vacuous bound that tracks data-complexity (see also Appendix E). This performance is not mirrored by several PAC–Bayes parameter‑space surrogates, which on the one hand are much less sensitive to training hyperparameters than the path norm measures are, but on the other hand exhibit a different more subtle kind of fragility; they are typically insensitive to label corruption in standard setups (see Appendix D). We call this behaviour a fragility because it reflects a serious failing of these measures.
@@
To illustrate data-complexity fragility, we chose the PACBAYES_ORIG measure (definition in App. A) because it is thought to be one of the best in its class. We hold the architecture and hyperparameters fixed and use the datasets MNIST, FashionMNIST, and CIFAR10. As illustrated in Figure 4, the dashed generalization–error curves separate cleanly and their slopes decrease with dataset difficulty (MNIST steepest decline, CIFAR‑10 flattest). However, the first two panels of Figure 4 show that for a given set of hyperparameters, the PACBAYES_ORIG measure gives nearly identical numerical values for different datasets and therefore fails to recognize changes in dataset complexity: its learning‑curve slopes are nearly indistinguishable across datasets and reflect changes in the test‑error slopes. By contrast, the third panel (ML‑PACBayes) is both tight (close to the corresponding error curves) and data‑aware: it preserves the MNIST << FashionMNIST << CIFAR‑10 ordering and reflects the different slopes. See Appendix E for the bound’s definition and a fuller discussion of why this function‑space quantity captures dataset difficulty while post-mortem parameter‑space PAC-Bayes surrogates often do not.
@@
In App. F, we move beyond label corruption and dataset swaps, and also probe symmetry‑preserving vs. signal‑destroying data transforms via pixel permutations. Again, we find that the ML-PACBayes function‑space predictor tracks the expected invariances while several post‑mortem diagnostics do not.
@@ 6 Scale-invariant network and exponential learning rate schedule
We use the same symmetry but a different equivalence. Instead of removing WD, our theorem pairs an exponentially increasing LR with a time‑varying WD so that every iterate computes the same predictor as a fixed‑LR/fixed‑WD baseline. The mapping is explicit and non‑asymptotic: it gives closed‑form schedules for $\tilde{\eta}_{t}$ and $\tilde{\lambda}_{t}$, and an admissible range for the multiplicative factor $\alpha$ via the interval $\mathcal{I}$. Conceptually, this creates a single “Exp++” knob that can drive large changes in parameter norms while leaving the learned function essentially unchanged—exactly the benign intervention we later use to probe the fragility of magnitude‑sensitive measures. Empirical details and results are deferred to Appendix G.
@@
Consequences. The proof of Theorem 6.2 is in Appendix C. In the scale‑invariant setting, this equivalence lets us pump parameter norms without changing $f(\theta)$, a lever we use to stress‑test generalization measures. Applying this dial, we find that even the measure that looked most stable in our earlier probes—the standard parameter‑space PAC–Bayes proxy PACBAYES_origfails catastrophically here: as the Exp++ factor grows, the proxy inflates by orders of magnitude while test error is essentially unchanged; see Appendix G.
@@ 7 Quantitative fragility: a scale‑free instability score
Both statistics are robust (medians) and scale‑free (invariant under $C\mapsto aC$, $a>0$). Unless noted otherwise we use absolute tolerances $\delta\in\{0.01,\,0.02,\,0.05\}$ (i.e., $1\%$ - $5\%$ of error), uniformly subsample pairs if they exceed a fixed budget, and report, for each measure, the medians $\mathrm{CMS}_{\delta}^{\mathrm{med}}$ and $\mathrm{eCMS}_{\delta}^{\mathrm{med}}$, along with the number of eligible pairs in (12)–(15) and the fraction of groups covered. Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters.
@@ 9 Alternative Views
It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalisation measures. This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (Soudry et al., 2018; Smith et al., 2017; Simsekli et al., 2019). From this perspective, the very invariances we prize risk washing out real, practically actionable differences between training recipes.
@@
A third objection is that using a marginal‑likelihood PAC–Bayes predictor as a baseline (Appendix E) overstates its practical authority. Since a GP is used to approximate the key marginal-likelihood term in the bound, the ML-PACBayes predictor could be seen as a measure of the performance of the underlying GP for any given architecture-data combination. While GPs are known to closely match DNN performance on the datasets we test in this paper (Lee et al., 2020b), it is hard to see how one could untangle the effects of hyperparameters or optimiser choice in this approach, in other words, the kinds of questions that post-mortem bounds are trying to address. More generally, modern deep-learning models are finite‑width, heavily augmented, and often trained far from the GP regime; priors that ignore augmentation, architectural quirks, or fine‑tuned tokenization may be badly mis‑specified. In addition, the bound controls the error of a stochastic Gibbs classifier drawn from a posterior over functions, not necessarily the deterministic network one actually deploys. Therefore, the GP route’s stability could stem from omitting factors that matter in practice (Lee et al., 2020b; Valle-Pérez & Louis, 2020).

### v2

@@ Abstract
A wide variety of generalization measures have been applied to deep neural networks (DNNs). Although obtaining tight bounds remains challenging, such measures are often assumed to reproduce qualitative generalization trends. In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks— are fragile: small training modifications that barely affect the underlying DNN can substantially change a measure’s value, trend, or scaling behavior. For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants can reverse the slope of a learning curve in widely used generalization measures like the path norm. We also identify subtler forms of fragility. For instance, the PAC-Bayes origin measure is regarded as one of the most reliable, and is indeed less sensitive to hyperparameter tweaks than many other measures. However, it completely fails to capture differences in data complexity across learning curves. This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behavior, in learning curves, but which is not a post-mortem measure. Beyond demonstrating that many bounds—such as path, spectral and Frobenius norms, flatness proxies, and deterministic PAC-Bayes surrogates—are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
@@ 1 Introduction
Classical generalization theories—based on VC dimension, Rademacher complexity and related tools—have a long history in machine learning. They were developed with two intertwined aims: to furnish tight mathematical guarantees and to explain generalization. For deep networks the same aspiration has driven a profusion of bounds. While tight bounds have proven elusive, with some exceptions (Dziugaite & Roy, 2017; Pérez‐Ortiz et al., 2020; Valle-Pérez & Louis, 2020; Lotfi et al., 2022), the expectation has been that these measures can generate qualitative insight into the generalization behavior of deep neural networks (DNNs). These bounds have typically focused on weights after training, which is why they are sometimes called post-mortem bounds. Large-scale comparisons have compared the behavior of these measures, albeit often only at the sign level, e.g. does changing a setting produce the same a change in the underlying DNN and a generalization measure with the same sign (Jiang et al., 2019a; Dziugaite et al., 2020a).
@@
The widespread fragility we observe suggests that some generalization measures may be showing correlation in some limited settings that may not be related to causation. Our position is therefore not only that many measures are fragile, but also that investigators should carefully audit their measures for fragility, using the kinds of tests we propose in this paper.
@@ 3 Training-hyperparameter fragility
In the left plot, we run SGD with momentum at $\mathrm{LR}=0.01$: both path‑norm curves start at a value of roughly ($\sim 10^{5}$) and decrease more or less log‑linearly with increasing sample size, with the margin‑normalized curve tracking about a decade lower. Both drop faster than the underlying ResNet‑50 test error. Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behavior. A minor adjjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows. These widely different behaviors arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
The contrasting behavior of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression. In that setting, for a fixed $\ell_{r}$ weight norm, qualitatively different scaling of the norm with sample size obtains depending on the $\ell_{p}$‑minimizer. More precisely, with $p\in[1,2]$, the $\ell_{r}$ weight norm is monotonic in $n$ if $r\leq 2(p-1)$ and plateaus if $r\leq 2(p-1)$  (Zhang & Louis, 2025). This implies that fixing which norm you measure does not fix its learning‑curve scaling, because the solution selected by optimization—through implicit (or explicit) bias, e.g., favoring different $\ell_{p}$‑minimizers—can induce different sample‑size scalings of that same norm. This change in qualitative behavior was also illustrated in diagonal linear networks where the implicit bias can be tuned by varying the learning dynamics (Zhang & Louis, 2025). For the more complex setup of path-norm for ResNet‑50 on FashinMNIST in Fig. 1, the optimizer and learning‑rate choices may analogously shift the deep model’s implicit bias, toggling the path‑norm scaling between monotone and U‑shaped, just as was seen for linear regression and deep linear networks.
@@ 5 Data-Complexity Fragility: Label Noise & Dataset Difficulty
Finally, the last panel in figure 3 illustrates how the function‑space ML‑PACBayes measure provides a non-vacuous bound that tracks data-complexity (see also Appendix F). This performance is not mirrored by several PAC–Bayes parameter‑space surrogates, which on the one hand are much less sensitive to training hyperparameters than the path norm measures are, but on the other hand exhibit a different more subtle kind of fragility; they are typically insensitive to label corruption in standard setups (see Appendix D). We call this behavior a fragility because it reflects a serious failing of these measures.
@@
To illustrate data-complexity fragility, we chose the PACBAYES_ORIG measure (definition in App. A) because it is thought to be one of the best in its class. We hold the architecture and hyperparameters fixed and use the datasets MNIST, FashionMNIST, and CIFAR10. As illustrated in Figure 4, the dashed generalization–error curves separate cleanly and their slopes decrease with dataset difficulty (MNIST steepest decline, CIFAR‑10 flattest). However, the first two panels of Figure 4 show that for a given set of hyperparameters, the PACBAYES_ORIG measure gives nearly identical numerical values for different datasets and therefore fails to recognize changes in dataset complexity: its learning‑curve slopes are nearly indistinguishable across datasets and reflect changes in the test‑error slopes. By contrast, the third panel (ML‑PACBayes) is both tight (close to the corresponding error curves) and data‑aware: it preserves the MNIST << FashionMNIST << CIFAR‑10 ordering and reflects the different slopes. See Appendix F for the bound’s definition and a fuller discussion of why this function‑space quantity captures dataset difficulty while post-mortem parameter‑space PAC-Bayes surrogates often do not.
@@
In App. G, we move beyond label corruption and dataset swaps, and also probe symmetry‑preserving vs. signal‑destroying data transforms via pixel permutations. Again, we find that the ML-PACBayes function‑space predictor tracks the expected invariances while several post‑mortem diagnostics do not.
@@ 6 Scale-invariant network and exponential learning rate schedule
We use the same symmetry but a different equivalence. Instead of removing WD, our theorem pairs an exponentially increasing LR with a time‑varying WD so that every iterate computes the same predictor as a fixed‑LR/fixed‑WD baseline. The mapping is explicit and non‑asymptotic: it gives closed‑form schedules for $\tilde{\eta}_{t}$ and $\tilde{\lambda}_{t}$, and an admissible range for the multiplicative factor $\alpha$ via the interval $\mathcal{I}$. Conceptually, this creates a single “Exp++” knob that can drive large changes in parameter norms while leaving the learned function essentially unchanged—exactly the benign intervention we later use to probe the fragility of magnitude‑sensitive measures. Empirical details and results are deferred to Appendix H.
@@
Consequences. The proof of Theorem 6.2 is in Appendix C. In the scale‑invariant setting, this equivalence lets us pump parameter norms without changing $f(\theta)$, a lever we use to stress‑test generalization measures. Applying this dial, we find that even the measure that looked most stable in our earlier probes—the standard parameter‑space PAC–Bayes proxy PACBAYES_origfails catastrophically here: as the Exp++ factor grows, the proxy inflates by orders of magnitude while test error is essentially unchanged; see Appendix H.
@@ 7 Quantitative fragility: a scale‑free instability score
Both statistics are robust (medians) and scale‑free (invariant under $C\mapsto aC$, $a>0$). Unless noted otherwise we use absolute tolerances $\delta\in\{0.01,\,0.02,\,0.05\}$ (i.e., $1\%$ - $5\%$ of error), uniformly subsample pairs if they exceed a fixed budget, and report, for each measure, the medians $\mathrm{CMS}_{\delta}^{\mathrm{med}}$ and $\mathrm{eCMS}_{\delta}^{\mathrm{med}}$, along with the number of eligible pairs in (12)–(15) and the fraction of groups covered. Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters. Full experimental details, formal algorithms and tabulated results can be found in Appendix E.
@@ 9 Alternative Views
It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalization measures. This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (Soudry et al., 2018; Smith et al., 2017; Simsekli et al., 2019). From this perspective, the very invariances we prize risk washing out real, practically actionable differences between training recipes.
@@
A third objection is that using a marginal‑likelihood PAC–Bayes predictor as a baseline (Appendix F) overstates its practical authority. Since a GP is used to approximate the key marginal-likelihood term in the bound, the ML-PACBayes predictor could be seen as a measure of the performance of the underlying GP for any given architecture-data combination. While GPs are known to closely match DNN performance on the datasets we test in this paper (Lee et al., 2020b), it is hard to see how one could untangle the effects of hyperparameters or optimiser choice in this approach, in other words, the kinds of questions that post-mortem bounds are trying to address. More generally, modern deep-learning models are finite‑width, heavily augmented, and often trained far from the GP regime; priors that ignore augmentation, architectural quirks, or fine‑tuned tokenization may be badly mis‑specified. In addition, the bound controls the error of a stochastic Gibbs classifier drawn from a posterior over functions, not necessarily the deterministic network one actually deploys. Therefore, the GP route’s stability could stem from omitting factors that matter in practice (Lee et al., 2020b; Valle-Pérez & Louis, 2020).